OpenStudy (anonymous):

A wholesaler of appliances finds that she can sell (2320 − p) flat-panel television sets each week when the price is p dollars. What price will maximize revenue? (Round your answer to the nearest cent.)

OpenStudy (dumbcow):

revenue = price*quantity R(p) = p*(2320-p) R(p) = 2320p -p^2 maximize by setting derivative to 0 R'(p) = 2320 - 2p = 0 2320 = 2p 1160 = p max R -> R(1160) = (2320*1160) - 1160^2 = 1,345,600